Robin is mowing a rectangular field that measures 24 yards by 32 yards, by pushing the mower around and around the outside of the plot. This creates a widening border that surrounds the unmowed grass in the center. During a brief rest, Robin wonders whether the job is half done yet. How wide is the uniform mowed border when Robin is half done?
Answer by checkley71(8403)

24*32=768 FT^2 IS THE TOTAL LAWN AREA.
(24-2X)(32-2X)=768/2
768-64X-48X+4X^2=384
4X^2-112X+768-384=0
4X^2-112X+384=0
4(X^2-28X+96)=0
4(X-24)(X-4)=0
X-4=0
X=4 ANSWER FOR THE WIDTH OF THE MOWED AREA.
(24-8)(32-8)=384
16*24=384
384=384